bill533:

Horse races are often measured in furlongs. If the Kentucky Derby is 10 furlongs long, what is its distance in miles? [1 furlong = 1/8 mile] A) 0.8 miles B) 1.25 miles C) 1.5 miles D) 4 miles E) 8 miles A platform has a volume of 135 cubic feet, but the concrete Joey wants to buy is sold by the cubic yard. How many cubic yards of concrete does Joey need to buy? A) 5 cubic yards B) 11.25 cubic yards C) 15 cubic yards D) 45 cubic yards E) 3645 cubic yards Michael's tennis serve is 105 mph. At the same rate, how many feet would his serve travel in one second to the nearest foot? A) 34 feet B) 75 feet C) 154 feet D) 260 feet E) 300 feet

bill533:

@Gdeinward

bill533:

1. C 2. B 3. E

bill533:

@Skatercat

Skatercat:

those are correct ^.^

bill533:

are they ?

bill533:

@Gdeinward

Gdeinward:

first one is B not C, 1 furlong = 1/8 mile 10 = 10/8 = 1 2/8 or 1.25

Gdeinward:

2nd one is D. 135 cubic feet, theres s feet in a yard, so divide the number of feet by 3 to get 45

Gdeinward:

the last one is definietley wrong, your answer should be smaller than the 105 miles

Gdeinward:

so A or B, but I dont know because that math is decently hard.

Gdeinward:

I know is 105 miles/ 1 hour so 105 miles to feet, then 1 to 60 seconds.

Gdeinward:

so 554400/ 60

Gdeinward:

then divide both by 60 to get ur answer.

bill533:

i's 1540

bill533:

it's

Gdeinward:

forget everyhting I just did.

bill533:

oh wait i divided 9240 divided by 6 instead of 60

bill533:

my bad

bill533:

it's 154

Gdeinward:

I dunno lemme ask dude to look at it, @dude

Gdeinward:

were experiencing technical difficulty on that last one.

dude:

Yeah you can convert miles to ft \(105mi\cdot \dfrac{5280ft}{1mi}=>554400ft\) Now from hours to [minutes to] seconds \(\dfrac{554400ft}{1h}\cdot \dfrac{1h}{60min}=>\dfrac{9240ft}{1min}\cdot\dfrac{1min}{60sec}=>154\) I think you just missed hour-> min-> seconds

Gdeinward:

yes ur right, I went from hours to seconds. Thanks Dude